OpenStudy (anonymous):

Hi Friends, I hope u can help me with this... Find the quadratic polynomial whose sum of zeros is -3 and product of zeros is 3 I am getting the answer as x^2 + 3x +3. But it doesn't seem right. Pls help....

OpenStudy (dumbcow):

if the zeroes are a and b then (x-a)(x-b) =0 a+b = -3 ab = 3 solve a and b using substitution a=-3-b (-3-b)b = 3 b^2 +3b+3=0 doesnt factor use quadratic formula b = -3+-sqrt(9-12)/2 b=-3/2 +-sqrt(3)/2 i plug back in to find a a=-3-(-3/2 +-sqrt(3)/2 i) a = -3/2 +-sqrt(3)/2 i so there is no real solution both zeroes are complex

OpenStudy (dumbcow):

i guess your answer would be correct y=x^2 +3x+3

OpenStudy (anonymous):

Thans "Dumbcow" for yr help.... A strange name u hv chosen!!

OpenStudy (dumbcow):

no prob, yeah i know it kinda an inside joke